OpenStudy (anonymous):

Solve and describe the steps you took to solve: 5m-7m+6=8m+2+10

OpenStudy (anonymous):

5m-7m+6=8m+2+10 Combine like terms: -2m+6=8m+12 Subtract 8m from both sides: -10m+6=12 Subtract 6 from both sides: -10m=6 Divide both sides by -10: m=-0.6

OpenStudy (rina.r):

yeah that's write i am donging it on notebook.

OpenStudy (rina.r):

*doing

OpenStudy (anonymous):

Ah, ok. Thanks very much!

OpenStudy (anonymous):

No prob

OpenStudy (rina.r):

Welcome:)